OpenStudy (anonymous):

if 18 grams of carbon react with oxygen to produce carbon dioxide, how many molecules of oxygen are required

OpenStudy (rajnie):

\[C + O2 \rightarrow CO2\] According to the balanced reaction 1 mole of Carbon atoms need 1 mole of Oxygen molecules. We have 18 grams of Carbon which is 1.5 moles of Carbon atoms and hence need 1.5 moles of Oxygen molecules. 1 mole of Oxygen molecules have Avogadro's Number of Oxygen Molecules. So, 1.5 moles of Oxygen will have 1.5 times Avogadro's number of molecules. The answer should be\[1.5 \times 6 \times 10^{23} = 9 \times 10^ {23} molecules of Oxygen\]

OpenStudy (anonymous):

thx rajnie